If 55 g of NaOH is added to enough distilled , deionized water to make a solution with a final volume of 1.25 L, what is the molarity of the resulting solution?
1
Expert's answer
2017-03-20T09:45:06-0400
C=n/V; n(NaOH)=m/M=55g/40g/Mol=1.375Mol; 1.375Mol/C=1.25L/1L;

C=1.375Mol/1.25L=1.1M